Two cars A and B are at town P. They want to travel to town Q. Car A starts 15 minutes earlier than car B and maintains a constant speed of 20 m/s whereas car B starts from rest and maintains a constant acceleration. Car A takes 30 minutes to reach the destination. If car B wants to reach the destination at the same time the acceleration required is 


1
Expert's answer
2021-08-03T11:42:32-0400

Given:

"t_A=30\:\rm min=1800\: s"

"t_B=15\:\rm min=900\: s"

"v_A=20\:\rm m/s=const"

"v_{Bi}=0"

The distance between Q and P

"d=v_A*t_A=1800*20=36000\:\rm m"


From the other side

"d=v_{Bi}t+a_Bt_B^2\/2=a_Bt_B^2\/2"


Hence, the acceleration of car B

"a_{Bf}=\frac{2d}{t_B^2}=\frac{2*36000}{900^2}=0.089\:\rm m/s^2"